OpenStudy (anonymous):

HELP ME The equation for finding the area of a triangle is A=1/2(b+c)h what is the equation solved for h.

OpenStudy (jhannybean):

to solve for h we're going to need to isolate h to one side of the equation. We have... \[\sf A= \frac{(b+c)h}{2}\] Let's start by distributing the h to both terms in the ()

OpenStudy (anonymous):

Mind if I work this out with you @Jhannybean or should I let this person try and work it out with you?

OpenStudy (jhannybean):

Sure, you can try

OpenStudy (anonymous):

So I would first start by multiplying both sides by two- A(2) = (b + c) h

OpenStudy (jhannybean):

Sure

OpenStudy (anonymous):

then I would subtract c from both sides- a(2) - c

OpenStudy (anonymous):

a(2) - c = b(h)

OpenStudy (jhannybean):

yeah, no. that doesnt work.

OpenStudy (jhannybean):

because now you have 2A = (b + c) \(\sf \color{red}{h}\) So you have to distribute h into all terms within the ( )

OpenStudy (anonymous):

but aren't we trying to figure out the formula for h?

OpenStudy (anonymous):

by distributing h, it would only make the problem more complex- yet you would get the same solution...right?

OpenStudy (jhannybean):

Well, yes, it would make it more complex. You are right. then what you can do is treat (b+c) as ONE term, and then divide both sides of the equation by it.

OpenStudy (anonymous):

Exactly! So a(2) - (b+c) = h correct?

OpenStudy (jhannybean):

\[\sf 2A=(b+c)h\]\[\sf h=\frac{2A}{b+c}\]

OpenStudy (anonymous):

Thank you guys